# SOLD: Reinforcement Learning with Slot Object-Centric Latent Dynamics

### v1

@@ B.2 Object-centric dynamics model
B.3 Slot Aggregation Transformer
@@ B.4 Training Details
Appendix C Baselines
In our experiments we compare our approach with two different baseline models, namely the SoTA model-based RL baseline DreamerV3 and a Non-Object-Centric variant of our proposed model:
@@ Appendix C Baselines
DreamerV3 (Hafner et al., 2023) is a SoTA model-based reinforcement learning algorithm that learns behaviors from visual inputs without requiring task-specific inductive biases or extensive environment interaction. It builds a world model that predicts future states and rewards, which is then used to simulate potential outcomes and guide the agent’s decision. DreamerV3 leverages latent dynamics and a compact, holistic representation of the environment for an efficient exploration, while showing desirable properties such as sample efficiency, scalability, and generalization across a wide range of complex tasks. We select the 12 million parameter variant so as to match the parameter count of our proposed model. For further details, we refer to Hafner et al. (2023).
@@
Appendix D Environments
In this section, we provide further details about our proposed suite of environments, which includes eight object-centric robotic control tasks designed to test relational reasoning and manipulation capabilities. The environments, which are inspired by Li et al. (2020) and are simulated using MuJoCo (Todorov et al., 2012), follow the same basic structure, consisting of a robot arm mounted on a base, positioned near a table where the manipulation tasks take place.
In all environments, the robot is controlled by a 4-dimensional action vector ${\bm{a}}=[a_{x},a_{y},a_{z},a_{grip}]\in[-1,1]^{4}$, where the first three components represent the desired movement direction of the end-effector, whereas the fourth component controls the gripper. On the Reach and Push tasks, commands to the gripper are ignored, with the gripper fixed in a closed configuration, as gripping is not required to solve these tasks.
@@ Appendix D Environments
In Reach tasks, the agent must identify a spherical target among several distractors and mode the end-effector to its location. The reward is calculated as:
@@
where ${\bm{p}}_{e}$ is the position of the end-effector and ${\bm{p}}_{t}$ is the target position. The success at the end of an episode is defined as:
@@
Push & PickAndPlace
Both Push and PickAndPlace correspond to reasoning and manipulation tasks where the agent must identify a single block among several distractors and move it to a target location. In Push tasks, the agent can slide the block to the target position on the table without using its gripper; whereas in PickAndPlace the target location can be above the table, thus needing to grasp the block in order to lift it to the target position. In both task variants the reward is calculated as:
@@
where ${\bm{p}}_{e}$ is the position of the end-effector, ${\bm{p}}_{t}$ is the target position, and ${\bm{p}}_{c}$ is the block position. The success at the end of an episode is defined as:
@@
Appendix E Additional Results
Figure 9 depicts the object-centric decomposition of a video frame obtained by SAVi. SAVi parses the input frame into per-object RGB reconstructions and alpha masks, which can be combined via a weighted sum in order to accurately reconstruct the observed video frame. Notably, SAVi assigns an object slot to the scene background, five slots to different blocks, one slot to the red target, and one slot to the robot arm. The sharp object masks demonstrate that SAVi isolates object-specific information in each slot, which is beneficial for downstream applications such as behavior learning, allowing the agent to reason about object properties and their relationships while abstracting task-irrelevant details.
@@ Appendix E Additional Results
We visualize action-conditional open-loop predictions on the Push Specific (Figure 10), Button-Press (Figure 11), Hammer (Figure 12), Cartpole-Balance (Figure 13), and Finger-Spin (Figure 14) environments. More precisely, we depict the ground truth sequence, the predicted video frames, the predicted instance segmentation of the scene, obtained by assigning a distinct color to each object mask, and the object reconstruction for each slot.
In all examples, our model parses the scene into sharp and accurate object representations, and models the action-conditional object dynamics and interactions in order to accurately predict the future video frames while preserving object-centric representations. We emphasize SOLD’s ability to capture complex physical interactions, such as pushing a block to a target location (10), pressing a button (11) or hitting a nail with a hammer (12). Furthermore, we showcase that SOLD can generalize to diverse non-object-centric environments (13 and 14).
@@
Figure 12: Open-loop prediction and decomposition results on the Hammer task. We visualize the ground truth and predicted video frames, instance segmentation obtained by assigning a different color to each object mask, and per-object reconstructions. SOLD assigns a slot for the scene background, three slots for different robot parts, a slot for the hammer, and a slot for the nail-box. Multiple slots are able to decompose the kinematic chain of the Sawyer robot in an unsupervised manner.
@@
Figure 16: Visualizing the actor’s attention over the slot-history on the PickAndPlace-Specific task reveals that the robot’s gripper and target block in the current time-step receive the most attention, highlighting the model’s ability to focus on task-relevant components. Further, the red target sphere is mostly occluded in the current time-step making it difficult to reconstruct its position accurately. However, the model learns to integrate information from previous steps, where it was still clearly visible.

### v2

@@ Appendix D Implementation Details
D.1 Slot Attention for Video
D.2 Object-centric dynamics model
D.3 Slot Aggregation Transformer
@@ D.4 Training Details
Appendix E Baselines
In our experiments we compare our approach with three different baseline models, namely the state-of-the-art model-based RL algorithms DreamerV3 and TD-MPC2 and a Non-Object-Centric variant of our proposed model.
@@ Appendix E Baselines
DreamerV3 (Hafner et al., 2023) is a SoTA model-based reinforcement learning algorithm that learns behaviors from visual inputs without requiring task-specific inductive biases or extensive environment interaction. It builds a world model that predicts future states and rewards, which is then used to simulate potential outcomes and guide the agent’s decision. DreamerV3 leverages latent dynamics and a compact, holistic representation of the environment for an efficient exploration, while showing desirable properties such as sample efficiency, scalability, and generalization across a wide range of complex tasks. We select the 12-million-parameter variant to match the parameter count of our proposed model. For further details, we refer to (Hafner et al., 2023).
TD-MPC2
TD-MPC2 (Hansen et al., 2024) builds upon its predecessor, TD-MPC (Hansen et al., 2022), with a series of architectural enhancements that improve scalability and sample efficiency. Like its predecessor, it avoids reconstructing high-dimensional inputs and instead focuses on modeling task-relevant dynamics in the latent space. The method employs temporal difference (TD) learning to predict future returns in the latent space and uses model predictive control (MPC) to optimize action sequences. Key advancements in TD-MPC2 include enhancements to training stability for larger model architectures and better generalization across tasks. These innovations allow it to achieve state-of-the-art performance on challenging visual and continuous control problems. For our experiments, we utilize the default 5-million-parameter variant since it is recommended by the authors for single-task RL problems.
@@
Appendix F Environments
In this section, we provide further details about our proposed suite of environments, which includes eight object-centric robotic control tasks designed to test relational reasoning and manipulation capabilities. The environments, which are inspired by (Li et al., 2020) and are simulated using MuJoCo (Todorov et al., 2012), follow the same basic structure, consisting of a robot arm mounted on a base, positioned near a table where the manipulation tasks take place.
In all environments, the robot is controlled by a 4-dimensional action vector ${\bm{a}}=[a_{x},a_{y},a_{z},a_{grip}]\in[-1,1]^{4}$, where the first three components represent the desired movement direction of the end-effector, whereas the fourth component controls the opening and closing of the gripper. On the Reach and Push tasks, commands to the gripper are ignored, with the gripper fixed in a closed configuration, as gripping is not required to solve these tasks.
@@ Appendix F Environments
In Reach tasks, the agent must identify a spherical target among several distractors and move the end-effector to its location. The reward is calculated as:
@@
where ${\bm{p}}_{e}$ is the position of the end-effector and ${\bm{p}}_{t}$ is the target position. Success is defined through the following condition at the last time step of an episode:
@@
Push & Pick
Both Push and Pick correspond to reasoning and manipulation tasks where the agent must identify a single block among several distractors and move it to a target location. In Push tasks, the agent can slide the block to the target position on the table without using its gripper; whereas in Pick the target location can be above the table, thus requiring the agent to grasp the block in order to lift it to the target position. In both task variants the reward is calculated as:
@@
where ${\bm{p}}_{e}$ is the position of the end-effector, ${\bm{p}}_{t}$ is the target position, and ${\bm{p}}_{c}$ is the block position. Success is defined through the following criterion, evaluated at the last time step of the episode:
@@
Appendix G Additional Results
G.1 Object-centric Decomposition
@@
Figure 8 depicts the object-centric decomposition of a video frame obtained by SAVi. SAVi parses the input frame into per-object RGB reconstructions and alpha masks, which can be combined via a weighted sum in order to accurately reconstruct the observed video frame. Notably, SAVi assigns an object slot to the scene background, five slots to different blocks, one slot to the red target, and one slot to the robot arm. The sharp object masks demonstrate that SAVi isolates object-specific information in each slot, which is beneficial for downstream applications such as behavior learning, allowing the agent to reason about object properties and their relationships while abstracting task-irrelevant details. Moreover, we find that SAVi is able to adapt to the varying number of objects in our environments, leaving extra slots empty when they are not needed to represent a scene.
G.2 Open-loop Prediction
We visualize action-conditional open-loop predictions in the Push-Specific (Figure 9), Button-Press (Figure 10), Hammer (Figure 11), Cartpole-Balance (Figure 12), and Finger-Spin (Figure 13) environments.
Specifically, we present the ground truth sequence, predicted video frames, instance segmentation – where each object mask is assigned a distinct color – and object reconstructions for each slot.
In all examples, our model parses the scene into sharp, accurate object representations and models action-conditional object dynamics and interactions, enabling precise future frame predictions while maintaining object-centric representations.
We highlight SOLD’s ability to capture complex physical interactions, such as pushing a block to a target location (Figure 9), pressing a button (Figure 10), or hammering a nail (Figure 11).
Furthermore, we demonstrate that SOLD generalizes to diverse, non-object-centric environments (Figure 12 and Figure 13), where sharp object separations and meaningful groupings emerge automatically – for instance, an object’s shadow, despite being spatially distinct, is assigned to the same slot as the object itself.
@@ G.2 Open-loop Prediction
Figure 11: Open-loop prediction and decomposition results on the Hammer task. We visualize the ground truth and predicted video frames, instance segmentation obtained by assigning a different color to each object mask, and per-object reconstructions. SOLD assigns a slot for the scene background, three slots for different robot parts, a slot for the hammer, and a slot for the nail-box.
@@
G.3 SAVi Fine-tuning
The Pick tasks highlight the need to design object-centric encoder-decoder modules that can adapt to changing state distribututions. Figure 14 exemplifies this challenge in the Specific variant, where the green target cube dissolves when lifted in the non-fine-tuned model.
Figure 15 underscores the importance of fine-tuning the object-centric encoder-decoder model with another example. Without fine-tuning, the blue color, which appears similarly on both colored blocks and the robot arm, leads to an even more drastic degradation of the reconstructions, where the robot itself is no longer accurately captured. In contrast, the fine-tuned model is able to reconstruct the sequence accurately.
True Observations
Frozen SAVi
Finetuned SAVi
Figure 15: Comparison of fine-tuned and frozen SAVi models on Pick-Distinct. We visualize the full reconstruction and the slot that reconstructs the cube that is being lifted for both models. When the blue block is lifted off the table, the frozen model merges it with blue elements from the robot arm, deteriorating the prediction and hallucinating the arm going between the gripper fingers. The fine-tuned model, on the other hand, is able to reconstruct the sequence accurately.
G.4 Discovering Task-relevant Objects
@@
Figure 17: Visualizing the actor’s attention over the slot-history on the Pick-Specific task reveals that the robot’s gripper and target block in the current time-step receive the most attention, highlighting the model’s ability to focus on task-relevant components. Further, the red target sphere is mostly occluded in the current time-step making it difficult to reconstruct its position accurately. However, the model learns to integrate information from previous steps, where it was still clearly visible.